### v1

@@ Abstract
Recent advances in foundation video generators such as Sora2, Veo3, and other commercial systems have produced highly realistic synthetic videos, exposing the limitations of existing detection methods that rely on shallow embedding trajectories, image-based adaptation, or computationally heavy MLLMs. We propose EA-Swin, an Embedding-Agnostic Swin Transformer that models spatiotemporal dependencies directly on pretrained video embeddings via a factorized windowed attention design, making it compatible with generic ViT-style patch-based encoders. Alongside the model, we construct the EA-Video dataset, a benchmark dataset comprising 130K videos that integrates newly collected samples with curated existing datasets, covering diverse commercial and open-source generators and including unseen-generator splits for rigorous cross-distribution evaluation. Extensive experiments show that EA-Swin achieves 0.97–0.99 accuracy across major generators, outperforming prior SoTA methods (typically 0.8–0.9) by a margin of 5–20%, while maintaining strong generalization to unseen distributions, establishing a scalable and robust solution for modern AI-generated video detection.
@@ 1 Introduction
Recent advances in generative artificial intelligence have led to a rapid transformation in video synthesis capabilities. Early video generation models in 2023 [chen2024videocrafter2, Khachatryan_2023_ICCV, wang2023modelscopetexttovideotechnicalreport] could only generate short, low-fidelity videos with limited temporal coherence. However, by 2025, hyper-realistic foundation models (e.g., Sora-2 [sora2] by OpenAI and Veo-3 [veo3] by Google) are capable of generating long, photorealistic videos from minimal input, including text prompts, reference images, or short video segments (see Figure 9). Powered by large-scale diffusion models [rombach2022high], transformers [vaswani2017attention, dosovitskiy2020image], and flow-matching [lipman2022flow] techniques, these systems can synthesize content that is increasingly difficult to distinguish from real-world footage, even defeating human perceptual detection capabilities in some cases. This has raised significant concerns about the use of Generative AI with malicious intentions, such as generation of inappropriate content and large-scale visual media fabrication [gambin2024deepfakes, YOON2025101491, easttom, CHEN2025108448].
@@
Although AI-generated video detection has garnered remarkable attention recently, existing approaches still exhibit notable limitations. These shortcomings primarily arise from two aspects: the detection methodologies themselves and the datasets used for evaluation. Some works [corvi2025seeing, zhang2025physicsdriven, interno2025aigenerated] propose detection strategies inspired by physics, mathematics, or geometric priors. While conceptually novel, subsequent experiments with more advanced generative models reveal that such handcrafted assumptions often fail to generalize, and in several cases, the reported performance gains can be largely attributed to strong pretrained vision backbones rather than the proposed mechanisms themselves. Other approaches [Zheng_2025_D3, yan2025orthogonal_effort, Chen_2025_forgelen], adapt image-based detection models to the video domain. However, this practice is fundamentally limited, as video generation introduces elements such as temporal dynamics, motion consistency, and long-range dependencies that cannot be captured by frame-level spatial analysis alone. Another line of methods relies on multimodal large language models (MLLMs) to distinguish real from synthetic videos in [wen2025busterxmllmpoweredaigeneratedvideo, park2025vidguardr1aigeneratedvideodetection, song2024on, li2025skyraaigeneratedvideodetection]. Despite their flexibility, these approaches are computationally expensive and impractical for large-scale or real-time deployment. Moreover, MLLMs primarily operate by producing semantic descriptions of visual content, rather than modeling the underlying generative process of videos, which limits their effectiveness as principled detection mechanisms. On the data side, existing benchmarks [genvidbench, wang2024vidprom], even when incorporating recently released generators, remain constrained by either outdated synthesis models or insufficient coverage of high-quality generators. In particular, videos from state-of-the-art commercial systems (e.g., Sora2 or Veo3) are often underrepresented due to high generation costs and restricted access, resulting in limited generator diversity and weakened generalization.
Our contributions in this paper are listed as follows:
@@
    We propose EA-Swin, a unified spatiotemporal architecture for AI-generated video detection. EA-Swin jointly models fine-grained spatial artifacts and temporal dynamics through an efficient window-based transformer design, while remaining compatible with generic ViT-style patch encoders.
@@
    We construct a large-scale benchmark dataset of 130K videos that combines newly collected samples with curated existing datasets, enabling comprehensive evaluation across diverse commercial and open-source generators, including unseen-generator splits for cross-distribution generalization.
@@
    We conduct extensive experiments across multiple vision backbones, which demonstrate that EA-Swin achieves SoTA accuracy and maintains a clear performance margin over recent competitive methods, validating its robustness for real-world AI-generated video detection.
@@ 2 Related Work
Recent years have seen a growing body of work on AI-generated video detection, moving beyond earlier studies that primarily focused on deepfake face manipulation or image-level synthetic content. Early works such as DeCoF [ma2025detectingaigeneratedvideoframe_decof] and DeMamba [demamba] represent some of the first attempts to explicitly address general AI-generated video detection, highlighting the need to model temporal artifacts beyond static visual cues. Existing approaches can be broadly categorized into video-based spatiotemporal models, embedding-trajectory–based methods, MLLM-based approaches, and image-based detectors commonly used for benchmarking.
@@
Embedding-trajectory-based methods analyze the temporal evolution of video representations extracted by pretrained encoders. Methods such as D3 measure simple differences between frame-level embeddings, while ResTraV [interno2025aigenerated_restrav] and NSG-VD [zhang2025physicsdriven_nsgvd] model higher-order temporal trajectories using statistics such as velocity, acceleration, or non-stationary graph structures; WaveRep [corvi2025waverep] further augments this paradigm by analyzing frequency-domain dynamics of embedding sequences. Despite their conceptual simplicity and efficiency, these methods face intrinsic limitations as video generators improve: embeddings from real and synthetic videos increasingly overlap in representation space, weakening trajectory separability. In particular, simpler methods like D3 [Zheng_2025_D3] become ineffective under modern generators, while ResTraV-style approaches (often relying on shallow MLP heads) lack sufficient capacity to capture deeper temporal dependencies, limiting their discriminative power and scalability.
More details on the AI-generated video detection method using image-based [univd23, fredect20, cnnspot20, gramnet20, Chen_2025_forgelen, yan2025orthogonal_effort, npr24] or MLLM-based [wen2025busterxmllmpoweredaigeneratedvideo, wen2026busterxunifiedcrossmodalaigenerated, li2025skyraaigeneratedvideodetection, park2025vidguardr1aigeneratedvideodetection, song2024_mm_det, fu2025learninghumanperceivedfakenessaigenerated, xiang2025aigvetoolaigeneratedvideoevaluation] methods are further discussed in the Supplementary.
@@ 2.2 Benchmark datasets
3 Method
Inspired by the Video Swin Transformer [liu2022videoswin], we propose a spatiotemporal detection framework named EA-Swin (Embedding-Agnostic Swin Transformer) that operates directly on pre-extracted video embeddings and captures their temporal and spatial dependencies using a Swin-style transformer head that factorizes temporal and spatial attention. Our design explicitly targets the temporal dynamics and spatial coherence of modern AI-generated videos, while remaining computationally efficient and encoder-agnostic.
3.1 Video Embedding Representation
Given a video $V$, we uniformly sample $T$ frames and extract visual representations using a frozen pretrained encoder. Depending on the backbone, embeddings may be frame-level or token-level. In the general case, each frame is decomposed into $S$ spatial tokens, each with embedding dimension of $D_{\text{in}}$. Thus, every video is represented as a 4D tensor $\mathbf{Z}\in\mathbb{R}^{B\times T\times S\times D_{\text{in}}}$, where $B$ denotes batch size. For frame-level encoders, we set $S=1$.
Optionally, consecutive frames may be grouped into temporal tubelets of size $\tau$ before projection, yielding $T^{\prime}=T/\tau$. For encoders without intrinsic temporal modeling (e.g., CLIP), we set $\tau=1$ to preserve frame-level dynamics. Instead of applying joint attention over all $T\times S$ tokens, we decompose the problem into alternating temporal and spatial attention blocks, substantially reducing computational complexity while preserving expressive power.
3.2 Factorized Spatiotemporal Swin Transformer
Figure 2: Spatiotemporal window shifting mechanism. The input video is first partitioned into non-overlapping local windows (1) along spatial and temporal dimensions. To enable cross-window interaction and enhance global context modeling, the windows are then shifted (2) spatially across adjacent regions and temporally across neighboring frames.
To enable effective interaction beyond local windows, we adopt a spatiotemporal window shifting strategy following the window-based attention mechanism. Specifically, the input video tokens are first partitioned into non-overlapping windows where self-attention is computed locally for efficiency. As shown in Figure 2, in the subsequent layer, the window partition is shifted along both spatial and temporal dimensions, allowing tokens near window boundaries to interact with neighboring regions and frames that were previously separated. This shifting operation improves information flow across windows, enhances spatiotemporal continuity, and enables the model to capture richer global context without significantly increasing computational cost.
Temporal Swin Attention.
We first model temporal dependencies independently for each spatial token (Figure 3). Specifically, we reshape $\mathbf{Z}$ as $\mathbf{Z}_{t}\in\mathbb{R}^{(B\cdot S)\times T\times D}$, treating each spatial token trajectory as an independent temporal sequence. Temporal attention is then applied using windowed multi-head self-attention with window size $W_{t}$. Let $\mathbf{z}\in\mathbb{R}^{W_{t}\times D}$ denote a temporal window.
Figure 3: Temporal Swin attention.
We compute
$\mathrm{Attn}_{t}(\mathbf{z})=$ (1)
$\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{h}}}+\mathbf{B}^{(t)}\right)V,$
where $Q,K,V\in\mathbb{R}^{W_{t}\times d_{h}}$ are query, key, and value projections for each head, $d_{h}=D/H$ is the per-head dimension, and $\mathbf{B}^{(t)}\in\mathbb{R}^{W_{t}\times W_{t}}$ is a learnable relative positional bias along the temporal axis.
Following the Swin Transformer design, alternating temporal blocks apply a cyclic shift of $W_{t}/2$ frames prior to window partitioning, enabling cross-window information flow while maintaining linear complexity in $T$. Each temporal block is followed by an MLP with residual connections and layer normalization.
Spatial Swin Attention.
After temporal modeling, we capture spatial interactions within each frame. If the number of spatial tokens satisfies $S=H_{p}\times W_{p}$, we reinterpret the tokens as a 2D grid and reshape into $\mathbf{Z}_{s}\in\mathbb{R}^{(B\cdot T)\times H_{p}\times W_{p}\times D}$.
Figure 4: Spatial Swin attention.
Then, we apply 2D windowed self-attention with window size $W_{s}\times W_{s}$. For each spatial window containing $N=W_{s}^{2}$ tokens, attention is computed as
$\mathrm{Attn}_{s}(\mathbf{z})=$ (2)
$\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{h}}}+\mathbf{B}^{(s)}\right)V,$
where $\mathbf{B}^{(s)}\in\mathbb{R}^{N\times N}$ is a learnable 2D relative positional bias encoding spatial offsets within the window.
As in the temporal case, spatial blocks alternate between non-shifted and shifted configurations (shift $W_{s}/2$ along both axes), enabling inter-window interaction while preserving locality (Figure 4). If the spatial tokens cannot be reshaped into a square grid, we fall back to 1D windowed attention along the token axis, ensuring robustness to arbitrary token layouts.
Figure 5: EA-Swin Architecture detail.
Block Composition.
The detection head consists of $D_{t}$ temporal Swin blocks followed by $D_{s}$ spatial Swin blocks (Figure 5). Each block follows the standard transformer structure:
$\mathbf{y}$ $=\mathbf{x}+\mathrm{MSA}(\mathrm{LN}(\mathbf{x})),$ (3)
$\mathbf{z}$ $=\mathbf{y}+\mathrm{MLP}(\mathrm{LN}(\mathbf{y})).$
This factorized design captures long-range temporal dynamics and fine-grained spatial coherence without incurring the quadratic cost of joint spatiotemporal attention.
3.3 Token Aggregation and Classification
After the final spatiotemporal block, all tokens are flattened to obtain $\mathbf{Z}_{\text{flat}}\in\mathbb{R}^{B\times(T\cdot S)\times D}$. We aggregate token features using either mean pooling or attention-based pooling, optionally incorporating a temporal validity mask. The resulting video-level representation is passed through a lightweight MLP classifier, which serves to predict whether the video is real or AI-generated.
@@ 4.1 Dataset Curation
In addition, we observe that many AI-generated videos are published on websites and social media platforms, making them a valuable data source. These videos are typically prompted by diverse users, are relatively long, and often undergo post-generation editing. To leverage this source, we scrape videos from websites offering AI video creation tools. According to their descriptions, these platforms generate videos using pretrained model and techniques such as prompt engineering, post-processing, or model fine-tuning. Specifically, we collect videos from OpenAI’s Sora [soraweb] website and other platforms, including DigenAI [digenai], ImaStudio [ima], Invideo [invideo], OpenArt [openart], and Pollo AI [pollo].
@@
After data collection, AI-generated videos are categorized by the generator. We then split the dataset into training, validation, and test sets. Specifically, generators with sufficient data (more than 3,0003{,}000 videos) are included in the training and validation sets, while generators with fewer samples are assigned to the test set, forming an unseen-generator benchmark. Real videos are split into training, validation, and test sets using the same ratios corresponding to each generator to ensure consistency across classes.
@@ 4.2 Dataset Composition
As shown in Figure 6 and 7, the final dataset comprises nearly 130K videos, including approximately 65K AI-generated videos and more than 62K real videos. The data are balanced across training, validation, and test splits for both real and AI videos, with comparable proportions in each split. The figure further illustrates the distribution of videos by generators and data sources.

### v2

@@ Abstract
Recent advances in foundation video generators such as Sora2, Veo3, and other commercial systems have produced highly realistic synthetic videos, exposing the limitations of existing detection methods that rely on shallow embedding trajectories, image-based adaptation, or computationally heavy MLLMs. We propose EA-Swin, an Embedding-Agnostic Swin Transformer that models spatiotemporal dependencies directly on pretrained video embeddings via a factorized windowed attention design, making it compatible with generic ViT-style patch-based encoders. Moreover, we construct the EA-Video dataset, a benchmark dataset comprising 130K videos that integrates newly collected samples with curated existing datasets, covering diverse commercial and open-source generators and including unseen-generator splits for rigorous cross-distribution evaluation. Extensive experiments show that EA-Swin achieves 0.97–0.99 accuracy across major generators, outperforming prior SoTA methods (typically 0.8–0.9) by a margin of 5–20%, while maintaining strong generalization to unseen distributions, establishing a scalable and robust solution for modern AI-generated video detection.
@@ 1 Introduction
Recent advances in generative artificial intelligence have led to a rapid transformation in video synthesis capabilities. Early video generation models in 2023 [chen2024videocrafter2, Khachatryan_2023_ICCV, wang2023modelscopetexttovideotechnicalreport] could only generate short, low-fidelity videos with limited temporal coherence. However, by 2025, hyper-realistic foundation models (e.g., Sora-2 [sora2] by OpenAI and Veo-3 [veo3] by Google) are capable of generating long, photorealistic videos from minimal input, including text prompts, reference images, or short video segments (see Figure 0.B.2). Powered by large-scale diffusion models [rombach2022high], transformers [vaswani2017attention, dosovitskiy2020image], and flow-matching [lipman2022flow] techniques, these systems can synthesize content that is increasingly difficult to distinguish from real-world footage, even defeating human perceptual detection capabilities in some cases. This has raised significant concerns about the use of Generative AI with malicious intentions, such as generation of inappropriate content and large-scale visual media fabrication [gambin2024deepfakes, YOON2025101491, easttom, CHEN2025108448].
@@
Although AI-generated video detection has attracted increasing attention, existing approaches still face fundamental limitations stemming from both methodology and data. Some works [corvi2025seeing, zhang2025physicsdriven, interno2025aigenerated] rely on physics-inspired or geometric priors. While conceptually appealing, these handcrafted assumptions often fail to generalize to modern high-fidelity generators, and their reported gains are frequently attributable to strong pretrained backbones rather than the proposed mechanisms. Other approaches [Zheng_2025_D3, yan2025orthogonal_effort, Chen_2025_forgelen] adapt image-based detectors to videos, which is inherently limited because video synthesis introduces temporal dynamics and long-range dependencies that cannot be captured by frame-level analysis alone. MLLM-based methods [wen2025busterxmllmpoweredaigeneratedvideo, park2025vidguardr1aigeneratedvideodetection, song2024on, li2025skyraaigeneratedvideodetection] offer flexibility but remain computationally expensive and unsuitable for large-scale deployment, while primarily relying on semantic reasoning rather than modeling the generative process itself. On the data side, existing benchmarks [genvidbench, wang2024vidprom] are often constrained by outdated generators or limited coverage of recent commercial models, leading to insufficient generator diversity and weak cross-distribution evaluation.
Meanwhile, rapid advances in foundation video generation have fundamentally changed the detection landscape. Early detectors operated in pixel space, where visible artifacts provided reliable forensic cues [ma2025detectingaigeneratedvideoframe_decof, corvi2025waverep]. Modern generators, however, are explicitly optimized to minimize pixel-level artifacts through diffusion models, transformers, and post-processing, making such cues increasingly unreliable. This shift suggests that detection must move beyond pixels and operate in representation space, where pretrained video encoders capture higher-level spatiotemporal dynamics that remain difficult for generative models to reproduce. While synthetic videos can achieve strong visual realism, they still struggle to match the temporal consistency and representation dynamics of real videos, motivating a representation-level paradigm for AI-generated video detection.
To explore this direction, we propose EA-Swin, an embedding-agnostic spatiotemporal detection head that operates directly on frozen video embeddings from foundation encoders. By decoupling detection from pixel-level processing, our approach enables scalable and robust detection for rapidly evolving video generators.
Our contributions are summarized as follows:
@@
    We introduce EA-Swin, an embedding-agnostic spatiotemporal detection framework that operates directly on frozen video representations, shifting AI-generated video detection from pixel space to representation space through a factorized Swin-style transformer that models temporal dynamics and spatial coherence in embedding space while remaining compatible with generic ViT-style encoders.
@@
    We construct EA-Video, a dataset of nearly 130K videos spanning commercial and open-source generators, and guarantee an unseen-generator protocol for cross-distribution evaluation.
@@
    Extensive experiments demonstrate consistent improvements over prior methods on both seen and unseen generators, validating representation-level spatiotemporal modeling as a robust solution for modern AI-generated video detection.
@@ 2 Related Work
Recent years have seen a growing body of work on AI-generated video detection, surpassing earlier studies that mainly focused on deepfake face manipulation or image-level synthetic content. Early works such as DeCoF [ma2025detectingaigeneratedvideoframe_decof] and DeMamba [demamba] represent some of the first attempts to explicitly address general AI-generated video detection, highlighting the need to model temporal artifacts beyond static visual cues. Existing approaches can be categorized into video-based spatiotemporal models, embedding-trajectory–based methods, MLLM-based approaches, and image-based detectors commonly used for benchmarking.
@@
Embedding-based methods analyze the temporal evolution of video representations extracted by pretrained encoders. Methods such as D3 measure simple differences between frame-level embeddings, while ResTraV [interno2025aigenerated_restrav] and NSG-VD [zhang2025physicsdriven_nsgvd] model higher-order temporal trajectories using statistics such as velocity, acceleration, or non-stationary graph structures; WaveRep [corvi2025waverep] further augments this paradigm by analyzing frequency-domain dynamics of embedding sequences. Despite their conceptual simplicity and efficiency, these methods face intrinsic limitations as video generators improve: embeddings from real and synthetic videos increasingly overlap in representation space, weakening trajectory separability. In particular, simpler methods like D3 [Zheng_2025_D3] become ineffective under modern generators, while ResTraV-style approaches (often relying on shallow MLP heads) lack sufficient capacity to capture deeper temporal dependencies, limiting their discriminative power and scalability.
More details on the AI-generated video detection method using image-based [univd23, fredect20, cnnspot20, gramnet20, Chen_2025_forgelen, yan2025orthogonal_effort, npr24] or MLLM-based [wen2025busterxmllmpoweredaigeneratedvideo, wen2026busterxunifiedcrossmodalaigenerated, li2025skyraaigeneratedvideodetection, park2025vidguardr1aigeneratedvideodetection, song2024_mm_det, fu2025learninghumanperceivedfakenessaigenerated, xiang2025aigvetoolaigeneratedvideoevaluation] methods are further discussed in the Supplementary Material.
@@ 3.1 Dataset Curation
In addition, we observe that many AI-generated videos are published on websites and social media platforms, making them a valuable data source. These videos are typically prompted by diverse users, are relatively long, and often undergo post-generation editing. To leverage this source, we collect AI-generated videos from publicly accessible platforms that provide video creation services. According to their descriptions, these platforms generate videos using pretrained models together with prompt engineering, post-processing, and fine-tuning strategies. We either create or obtain videos through OpenAI’s Sora [soraweb] and other platforms, including DigenAI [digenai], ImaStudio [ima], Invideo [invideo], OpenArt [openart], and Pollo AI [pollo].
@@
After data collection, AI-generated videos are categorized by the generator. We then split the dataset into training, validation, and test sets. Specifically, generators with sufficient data (more than 3,0003{,}000 videos) are included in the training and validation sets, while generators with fewer samples are assigned to the test set, forming an unseen-generator benchmark. Real videos are split into training, validation, and test sets using the same ratios corresponding to each generator to ensure consistency across classes. More details about the dataset can be found in Supplementary Material.
@@ 3.2 Dataset Composition
As shown in Figure 2 and 3, the final dataset comprises 127,054 videos, including approximately 65K AI-generated videos and more than 62K real videos. The data are balanced across training, validation, and test splits for both real and AI videos, with comparable proportions in each split. The figure further illustrates the distribution of videos by generators and data sources.
@@
4 Method
4.1 Representation Trajectory Analysis
To understand how real and AI-generated videos differ in representation space, we project frame-level embeddings from a pretrained video encoder into 2D using t-SNE and visualize their temporal trajectories (8 frames per video). Each polyline corresponds to the embedding evolution of a single video across time.
Figure 4: t-SNE visualization of embedding trajectories. Each polyline represents the temporal evolution of video embeddings from V-JEPA 2 encoder.
As shown in Figure 4, real and AI-generated videos partially overlap at early frames but gradually diverge as temporal dynamics unfold. While real videos exhibit diverse and irregular trajectory patterns, AI-generated videos tend to drift toward more concentrated regions with smoother and more constrained transitions. This suggests that although modern generators can closely match pixel-level appearance, they fail to fully reproduce the spatiotemporal dynamics captured by pretrained video representations.
These observations indicate that temporal evolution in embedding space provides a stronger forensic signal than static frame-level analysis, motivating a detection framework that explicitly models representation trajectories rather than raw pixels.
4.2 Embedding-Agnostic Spatiotemporal Modeling
Based on the above analysis, we design a lightweight spatiotemporal detection head that operates directly on frozen video embeddings. Unlike Video Swin [liu2022videoswin], which processes high-dimensional pixel inputs using large spatial windows as a full video backbone, our setting operates on compact pretrained embeddings that already encode rich semantic and motion information. Therefore, instead of re-learning visual representations from pixels, we focus on modeling the temporal evolution and spatial coherence of embedding trajectories.
To this end, we propose EA-Swin (Embedding-Agnostic Swin Transformer), a factorized Swin-style transformer that performs temporal and spatial attention in embedding space. By decoupling detection from pixel-level processing, EA-Swin remains computationally efficient, encoder-agnostic, and specifically tailored for AI-generated video forensics.
Figure 5: Spatiotemporal window shifting mechanism. The input video embedding is first partitioned into non-overlapping local windows (1) along spatial and temporal dimensions. To enable cross-window interaction and enhance global context modeling, the windows are then shifted (2) spatially across adjacent regions and temporally across neighboring frames.
4.3 Embedding-Agnostic Spatiotemporal Detection Head
Video embedding representation.
Given a video $V$, we uniformly sample $T$ frames and extract features using a frozen pretrained video encoder. Depending on the backbone, embeddings may be frame-level or token-level. In the general case, each frame is decomposed into $S$ spatial tokens with dimension $D_{\text{in}}$, yielding a 4D representation
$\mathbf{Z}\in\mathbb{R}^{B\times T\times S\times D_{\text{in}}},$
where $B$ denotes the batch size. For frame-level encoders we set $S=1$. Since pretrained encoders already capture rich semantic and motion information, our goal is not to relearn visual features from pixels but to model the spatiotemporal evolution of embedding trajectories.
Figure 6: Temporal Swin attention.
Factorized spatiotemporal Swin attention.
To model representation dynamics efficiently, we design a lightweight Swin-style detection head that alternates temporal and spatial windowed attention (Fig. 5). Instead of applying joint attention over all $T\times S$ tokens, this factorized design significantly reduces computational cost while preserving long-range modeling capability. The window shifting mechanism enables information exchange across neighboring frames and spatial regions without incurring quadratic complexity.
Temporal Swin attention.
Figure 7: Spatial Swin attention.
We first model temporal dependencies independently for each spatial token (Fig. 6) by reshaping
$\mathbf{Z}_{t}\in\mathbb{R}^{(B\cdot S)\times T\times D}.$
Windowed multi-head self-attention with window size $W_{t}$ is applied:
$\mathrm{Attn}_{t}(\mathbf{z})=\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{h}}}+\mathbf{B}^{(t)}\right)V,$
where $\mathbf{B}^{(t)}$ is a learnable temporal relative positional bias. Alternating shifted windows allow cross-frame interaction while maintaining linear complexity in $T$.
Spatial Swin attention.
After temporal modeling, spatial interactions within each frame are captured by reshaping tokens into a grid
$\mathbf{Z}_{s}\in\mathbb{R}^{(B\cdot T)\times H_{p}\times W_{p}\times D}.$
We then apply 2D windowed attention
$\mathrm{Attn}_{s}(\mathbf{z})=\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{h}}}+\mathbf{B}^{(s)}\right)V,$
where $\mathbf{B}^{(s)}$ encodes spatial relative positional bias (Fig. 7). Shifted windows again enable inter-window communication while preserving locality.
Figure 8: EA-Swin architecture.
Detection head and classification.
The detection head consists of $D_{t}$ temporal blocks followed by $D_{s}$ spatial blocks (Fig. 8). Each block follows the standard transformer formulation
$\mathbf{y}=\mathbf{x}+\mathrm{MSA}(\mathrm{LN}(\mathbf{x})),\quad\mathbf{z}=\mathbf{y}+\mathrm{MLP}(\mathrm{LN}(\mathbf{y})).$
After the final block, tokens are flattened and pooled to obtain a video-level representation, which is fed into a lightweight MLP classifier to predict whether the video is real or AI-generated.